# Scaling Beyond Context: A Survey of Multimodal Retrieval-Augmented Generation for Document Understanding

### v1

@@ 4 Dataset and Benchmark
DocVQA (2021) 50K 12,767 (I) q !
VisualMRC (2021) 30,562 10,197 (I) q !
TAT-DQA (2022) 16,558 2,758 (D) q !
InfoVQA (2022) 30K 5.4K (I) q !
@@
ScienceQA (2022) 21K 7,803 (I) q !
DUDE (2023) 41,491 4,974 (D) q !
SlideVQA (2023) 52K 14.5K (I) 1
ArXivQA (2024a) 100K 16.6K (D) q !
MMLongBench-Doc (2024b) 1,062 130 (D) q !  1
PaperTab (2024) 393 307 (D) q !
FetaTab (2024) 1,023 878 (D) !
SPIQA (2024) 27K 25.5K (D) !
LongDocUrl (2024) 2,325 396 (D) q !
ViDoRe (2024) 3.8K 8.3K (D) q !
VisR-Bench (2024b) 471 226 (D) q !  1
M3DoCVQA (2024a) 2,441 3,368 (D) q !
VisDoMBench (2025) 2,271 1,277 (D) q !  1
ViDoSeek (2025b) 1,142 300 (D) q !
OpenDocVQA (2025) 206K 43K (I) q !
Table 2: Overview of datasets and benchmarks in multimodal RAG for document understanding. We report the number of queries, dataset size, and covered content types (q Text, ! Tables, Charts, 1 Slides). (D) and (I) indicate that the count refers to documents or images, respectively. The upper part covers widely used multimodal document understanding datasets; the lower part compiles recent multimodal RAG benchmarks introduced by methods surveyed in this paper to address prior limitations.
@@
Many approaches have also identified various shortcomings in multimodal RAG, which has led to the construction of diverse benchmarks. These newly proposed benchmarks are summarized in the lower part of Table 2. ColPali (Faysse et al., 2024) constructs a comprehensive and diverse benchmark, ViDoRe, which includes both academic and practical tasks across domains such as energy, government, and healthcare, while SV-RAG (Chen et al., 2024b) collects a human-verified dataset of great diversity to establish VISR-BENCH. Similarly, M3DocVQA (Cho et al., 2024a), VisDoMRAG (Suri et al., 2025) and VDocRAG (Tanaka et al., 2025) address the limitation that most existing benchmarks focus on single-document retrieval rather than cross-document open-domain scenarios by introducing the open-domain benchmarks M3DocVQA, VisDoMBench and OpenDocVQA, respectively. To further overcome the constraint that existing VQA datasets typically pair queries with only one or a few images, whereas real-world applications require reasoning over large-scale corpora, ViDoRAG (Wang et al., 2025b) introduces ViDoSeek, a novel visually rich document dataset specifically tailored for RAG systems, enabling queries with unique answers in realistic retrieval settings and supporting rigorous evaluation of multimodal document understanding.
We also present the performance of different multimodal RAG methods across various benchmarks, along with a detailed explanation of the evaluation metrics and their computation. The full details are provided in the appendix.
@@ 5 Application
6 Challenge and Future Direction
Despite continued advances in multimodal RAG for document understanding, several open challenges direct future research. First, efficiency is a central concern (Ma et al., 2025): integrating high-dimensional visual and textual features incurs substantial computational cost, constraining scalability. Promising directions include lightweight multimodal encoders, adaptive retrieval, and memory-efficient fusion that reduce latency without sacrificing retrieval accuracy. Second, finer-grained document representations are necessary (Wang et al., 2025c; Xu et al., 2025b; Yu et al., 2025a; Gong et al., 2025; Choi et al., 2025): many models operate at the page or paragraph level and overlook tables, figures, footnotes, and layout-specific semantics. Hierarchical encoders and attention mechanisms that capture micro-level structure while preserving collection-level context can improve interpretability and enhance downstream reasoning and decision-making. Finally, security and robustness (Shereen et al., 2025; Cho et al., 2024b; Nazary et al., 2025; Jiang et al., 2024; Xian et al., 2024) rise in importance as adoption expands in finance, healthcare, and law: systems must resist adversarial manipulation, data leakage, and hallucination risk. Privacy-preserving multimodal retrieval, verifiable generation, and trust-calibration mechanisms are essential to ensure reliability and regulatory compliance.
@@ 7 Conclusion
Although this survey aims to provide a comprehensive synthesis of multimodal RAG for document understanding, several limitations remain. First, while we highlight practical applications, our analysis of real-world deployment challenges such as industry adoption, user-centered evaluation, system integration, and deployment scalability remains preliminary. Broader socio-technical aspects of multimodal RAG systems deserve further exploration in future work. Second, although we summarize major datasets and benchmarks, a more systematic investigation into data quality, annotation consistency, inter-domain transferability, and evaluation alignment across modalities would provide deeper insights into their generalizability and real-world relevance. Furthermore, as multimodal RAG for document understanding is an emerging and rapidly evolving field, newly released datasets, models, and evaluation protocols continue to reshape the landscape. To address this dynamic nature, this survey will be periodically updated and complemented by an open repository to track ongoing progress and facilitate community collaboration.
@@ Appendix A Dataset and Benchmark
DSE (Ma et al., 2024a) R@10 84.6
VisRAG (Yu et al., 2024) R@10 91.20 97.39 97.08
CMRAG (Chen et al., 2025) R@10 64.12
VisRAG (Yu et al., 2024) MRR@10 75.37 91.85 86.37
CMRAG (Chen et al., 2025) MRR@10 47.64
ColPali (Faysse et al., 2024) nDCG@5 54.4 81.8
ColQwen2 (Faysse et al., 2024) nDCG@5 61.5 89.7
ColQwen2.5 (Faysse et al., 2024) nDCG@5 63.6 92.5
VDocRAG (Tanaka et al., 2025) nDCG@5 77.3 72.9
@@
DSE (Ma et al., 2024a) nDCG@10 75.3
CMRAG (Chen et al., 2025) nDCG@10 52.10
@@
VisRAG (Yu et al., 2024) EM 67.17 60.97 66.43
FRAG (Huang et al., 2025) EM 72.7
SV-RAG (Chen et al., 2024b) PNLS 87.0 98.8 84.8
CRAEM (Zhang et al., 2024a) ANLS 79.4 53.6
M3DocRAG (Cho et al., 2024a) ANLS 84.4
VisDoMRAG (Suri et al., 2025) ANLS 67.2
VDocRAG (Tanaka et al., 2025) ANLS 56.4 64.6
FRAG (Huang et al., 2025) ANLS 87.4
ReDocRAG (López et al., 2025) ANLS 73.7 63.6
M3DocRAG (Cho et al., 2024a) G-Acc 21.0
FRAG (Huang et al., 2025) G-Acc 80.5 37.9
VRAG-RL (Wang et al., 2025c) G-Acc 69.3 24.9
SimpleDoc (Jain et al., 2025) G-Acc 60.58
MMRAG-DocQA (Gong et al., 2025) G-Acc 52.3
CMRAG (Chen et al., 2025) G-Acc 43.25
MoLoRAG (Wu et al., 2025) G-Acc 41.01
@@
In addition, we compile the evaluation results of various multimodal RAG methods on widely used benchmarks, including DocVQA (Mathew et al., 2021), InfoVQA (Mathew et al., 2022), SlideVQA (Tanaka et al., 2023), and MMLongBench-Doc (Ma et al., 2024b), as presented in Table A. These results provide a clear comparison of the strengths and weaknesses of different approaches. The evaluation of multimodal RAG performance typically falls into two categories: retrieval and generation, which are presented in the upper and lower parts of Table A, respectively. Retrieval evaluation focuses on the accuracy of the retrieved pages, whereas generation evaluation measures the correctness of model outputs when the retrieved pages are combined with the user query as input. Since different methods adopt slightly different metrics, we annotate these variations in the table, while aligning comparable metrics to facilitate direct comparison. Detailed explanations of these metrics are provided in Appendix Section B.
@@ B.1 Retrieval Evaluation
The F1-Score is often adopted as the harmonic mean of Precision@K and Recall@K, widely used to assess the correctness of retrieved entities or factual fragments in both the retrieval module and the generation process (Li et al., 2024b):
@@ B.2 Generation Evaluation
In the context of Multimodal RAG, the primary objective of generation quality evaluation is to assess the quality and consistency between model-generated text and reference answers. This involves not only measuring the correctness of the responses but also considering aspects such as fluency, information coverage, and logical coherence. To achieve a comprehensive evaluation, this study examines a wide range of metrics. The earliest are soft matching metrics (e.g., BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005)), which rely on n-gram overlap for a soft lexical evaluation that allows partial and flexible matching. They mainly assess fluency and information coverage of generated text. With the rise of question answering and reading comprehension tasks, strict matching metrics (e.g., Exact Match (Rajpurkar et al., 2016), ANLS (Biten et al., 2019), PNLS (Chen et al., 2024a)) are introduced, focusing on exact or near-exact correspondence with reference answers to measure form-level correctness. More recently, driven by the advances in pretrained language models, semantic matching metrics (e.g., BERTScore (Devlin et al., 2019), RoBERTa (Liu et al., 2019), G-Acc (Ma et al., 2024b)) have become prominent, enabling the assessment of deeper semantic consistency through contextual embeddings. By combining these three categories of metrics, generation quality can be evaluated holistically across surface, exact matching, and semantic alignment.
@@
Beyond soft and strict matching metrics, semantic matching metrics have emerged to evaluate deeper semantic consistency between generated and reference texts. Metrics such as BERTScore, which leverages contextual embeddings from pretrained language models like BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), move beyond simple lexical overlap by capturing semantic similarity between generated and reference texts. This enables a more reliable evaluation of whether the meaning of a response is preserved, even when different phrasings are used. However, while BERTScore provides strong advantages in measuring semantic consistency, it is less suited for scenarios involving long-form, explanatory, or unanswerable responses. To address this gap, Generated Accuracy (G-Acc) (Ma et al., 2024b) has been proposed, which extends evaluation to free-form answers that emphasize reasoning, elaboration, and contextual completeness, thereby offering a more comprehensive assessment of generation quality.

### v2

@@ 4 Dataset and Benchmark
DocVQA (2021) 50K 12,767 (I)
VisualMRC (2021) 30,562 10,197 (I)
TAT-DQA (2022) 16,558 2,758 (D)
InfoVQA (2022) 30K 5.4K (I)
@@
ScienceQA (2022) 21K 7,803 (I)
DUDE (2023) 41,491 4,974 (D)
SlideVQA (2023) 52K 14.5K (I)
ArXivQA (2024a) 100K 16.6K (D)
MMLongBench-Doc (2024c) 1,062 130 (D)
PaperTab (2024) 393 307 (D)
FetaTab (2024) 1,023 878 (D)
SPIQA (2024) 27K 25.5K (D)
LongDocUrl (2024a) 2,325 396 (D)
ViDoRe (2024) 3.8K 8.3K (D)
VisR-Bench (2024b) 471 226 (D)
M3DoCVQA (2024a) 2,441 3,368 (D)
VisDoMBench (2025) 2,271 1,277 (D)
ViDoSeek (2025b) 1,142 300 (D)
OpenDocVQA (2025) 206K 43K (I)
UniDoc-Bench (2025) 1.6K 70K(I)
BBox-DocVQA (2025b) 32K 4.4K(D)
Table 3: Overview of datasets and benchmarks in multimodal RAG for document understanding. We report the number of queries, dataset size, and covered content types ( Text, Tables, Charts, Slides). (D) and (I) indicate that the count refers to documents or images, respectively. The upper part covers widely used multimodal document understanding datasets; the lower part compiles recent multimodal RAG benchmarks introduced by methods surveyed in this paper to address prior limitations.
@@
Many studies have revealed limitations in multimodal RAG systems, leading to the development of diverse benchmarks summarized in the lower half of Table 3. ColPali (Faysse et al., 2024) proposes ViDoRe, a comprehensive benchmark covering academic and practical tasks across domains such as energy, government, and healthcare; while SV-RAG (Chen et al., 2024b) builds VISR-BENCH from a large-scale, manually validated dataset with high task diversity. To overcome single-document evaluation, M3DocVQA (Cho et al., 2024a), VisDoMRAG (Suri et al., 2025), and VDocRAG (Tanaka et al., 2025) extend evaluation to cross-document open-domain scenarios using M3DocVQA, VisDoMBench, and OpenDocVQA, respectively. Focusing on large-scale retrieval closer to real-world applications, ViDoRAG (Wang et al., 2025b) introduces ViDoSeek, a benchmark for RAG evaluation with uniquely answerable queries. Furthermore, UniDoc-Bench (Peng et al., 2025) establishes a document-centric MM-RAG benchmark, enabling systematic comparisons of multimodal retrieval and fusion strategies on real-world PDFs through unified text, table, and figure evidence linking. BBox-DocVQA (Yu et al., 2025b) provides a DocVQA dataset with bounding-box grounding for supervision of spatial reasoning and evidence localization.
We also present the performance of different multimodal RAG methods across various benchmarks, along with a detailed explanation of the evaluation metrics and their computation. The full details are provided in Appendix A.
@@ 5 Application
6 Challenge, Critical Analysis and Industry Deployment
Due to space constraints, extended discussions are deferred to the appendix. Appendix D outlines key open challenges and future directions in multimodal RAG, focusing on efficiency, training paradigms, granularity, and security. Appendix E presents a concise critical analysis of fundamental limitations and representative failure cases beyond aggregate benchmarks. Appendix F addresses industrial deployment considerations, highlighting practical constraints, efficiency trade-offs, and representative open-source systems.
@@ 7 Conclusion
Although this survey aims to provide a comprehensive synthesis of multimodal RAG for document understanding, several limitations remain. First, while we highlight practical applications, our analysis of real-world deployment challenges such as user-centered evaluation, system integration, and deployment scalability remains preliminary. Broader socio-technical aspects of multimodal RAG systems deserve further exploration in future work. Second, although we summarize major datasets and benchmarks, a more systematic investigation into data quality, annotation consistency, inter-domain transferability, and evaluation alignment across modalities would provide deeper insights into their generalizability and real-world relevance. Furthermore, as multimodal RAG for document understanding is an emerging and rapidly evolving field, newly released datasets, models, and evaluation protocols continue to reshape the landscape. To address this dynamic nature, this survey will be periodically updated and complemented by an open repository to track ongoing progress and facilitate community collaboration.
@@ Appendix A Dataset and Benchmark
   SV-RAG (Chen et al., 2024b)    Top-5    87.0    98.8    –    84.8
   DSE (Ma et al., 2024b)    R@10    –    84.6    –    –
   VisRAG (Yu et al., 2024)    R@10    91.20    97.39    97.08    –
   CMRAG (Chen et al., 2025c)    R@10    –    –    –    64.12
   RegionRAG (Li et al., 2025b)    R@10    99.4    98.4    99.5    –
   VisRAG (Yu et al., 2024)    MRR@10    75.37    91.85    86.37    –
   CMRAG (Chen et al., 2025c)    MRR@10    –    –    –    47.64
   LILaC (Yun et al., 2025)    MRR@10    78.75    84.43    86.83    –
   ColPali (Faysse et al., 2024)    nDCG@5    54.4    –    81.8    –
   ColQwen2 (Faysse et al., 2024)    nDCG@5    61.5    –    89.7    –
   ColQwen2.5 (Faysse et al., 2024)    nDCG@5    63.6    –    92.5    –
   VDocRAG (Tanaka et al., 2025)    nDCG@5    –    77.3    72.9    –
@@
   RegionRAG (Li et al., 2025b)    nDCG@5    93.1    90.3    94.8    –
   HKRAG (Tong et al., 2025)    nDCG@5    –    74.3    71.9    –
   DSE (Ma et al., 2024b)    nDCG@10    –    75.3    –    –
   CMRAG (Chen et al., 2025c)    nDCG@10    –    –    –    52.10
@@
   VisRAG (Yu et al., 2024)    EM    67.17    60.97    66.43    –
   FRAG (Huang et al., 2025a)    EM    –    72.7    –    –
   LILaC (Yun et al., 2025)    EM    65.48    55.57    60.91    –
   SV-RAG (Chen et al., 2024b)    PNLS    76.0    77.0    –    49.0
   CRAEM (Zhang et al., 2024a)    ANLS    79.4    –    53.6    –
   M3DocRAG (Cho et al., 2024a)    ANLS    84.4    –    –    –
   VisDoMRAG (Suri et al., 2025)    ANLS    –    67.2    –    –
   VDocRAG (Tanaka et al., 2025)    ANLS    –    56.4    64.6    –
   FRAG (Huang et al., 2025a)    ANLS    87.4    –    –    –
   ReDocRAG (López et al., 2025)    ANLS    73.7    –    63.6    –
   M3DocRAG (Cho et al., 2024a)    G-Acc    –    –    –    21.0
   FRAG (Huang et al., 2025a)    G-Acc    80.5    –    –    37.9
   VRAG-RL (Wang et al., 2025c)    G-Acc    –    69.3    –    24.9
   SimpleDoc (Jain et al., 2025)    G-Acc    –    –    –    60.58
   MMRAG-DocQA (Gong et al., 2025)    G-Acc    –    –    –    52.3
   CMRAG (Chen et al., 2025c)    G-Acc    –    –    –    43.25
   MoLoRAG (Wu et al., 2025b)    G-Acc    –    –    –    41.01
   RECON (Wang and Chen, 2025)    G-Acc    –    66.12    –    –
   LAD-RAG (Sourati et al., 2025)    G-Acc    82.9    –    –    45.0
   DREAM (Zhang et al., 2025a)    G-Acc    –    –    –    27.3
   MARA (Wu et al., 2025a)    G-Acc    84.64    73.40    68.02    –
   SLEUTH (Liu et al., 2025a)    G-Acc    –    –    –    52.77
@@
In addition, we compile the evaluation results of various multimodal RAG methods on widely used benchmarks, including DocVQA (Mathew et al., 2021), InfoVQA (Mathew et al., 2022), SlideVQA (Tanaka et al., 2023), and MMLongBench-Doc (Ma et al., 2024c), as presented in Table A. These results provide a clear comparison of the strengths and weaknesses of different approaches. The evaluation of multimodal RAG performance typically falls into two categories: retrieval and generation, which are presented in the upper and lower parts of Table A, respectively. Retrieval evaluation focuses on the accuracy of the retrieved pages, whereas generation evaluation measures the correctness of model outputs when the retrieved pages are combined with the user query as input. Since different methods adopt slightly different metrics, we annotate these variations in the table, while aligning comparable metrics to facilitate direct comparison. Detailed explanations of these metrics are provided in Appendix B.
@@ B.1 Retrieval Evaluation
The F1-Score is often adopted as the harmonic mean of Precision@K and Recall@K, widely used to assess the correctness of retrieved entities or factual fragments in both the retrieval module and the generation process (Li et al., 2024c):
@@ B.2 Generation Evaluation
In the context of Multimodal RAG, the primary objective of generation quality evaluation is to assess the quality and consistency between model-generated text and reference answers. This involves not only measuring the correctness of the responses but also considering aspects such as fluency, information coverage, and logical coherence. To achieve a comprehensive evaluation, this study examines a wide range of metrics. The earliest are soft matching metrics (e.g., BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005)), which rely on n-gram overlap for a soft lexical evaluation that allows partial and flexible matching. They mainly assess fluency and information coverage of generated text. With the rise of question answering and reading comprehension tasks, strict matching metrics (e.g., Exact Match (Rajpurkar et al., 2016), ANLS (Biten et al., 2019), PNLS (Chen et al., 2024a)) are introduced, focusing on exact or near-exact correspondence with reference answers to measure form-level correctness. More recently, driven by the advances in pretrained language models, semantic matching metrics (e.g., BERTScore (Devlin et al., 2019), RoBERTa (Liu et al., 2019), G-Acc (Ma et al., 2024c)) have become prominent, enabling the assessment of deeper semantic consistency through contextual embeddings. By combining these three categories of metrics, generation quality can be evaluated holistically across surface, exact matching, and semantic alignment.
@@
Beyond soft and strict matching metrics, semantic matching metrics have emerged to evaluate deeper semantic consistency between generated and reference texts. Metrics such as BERTScore, which leverages contextual embeddings from pretrained language models like BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), move beyond simple lexical overlap by capturing semantic similarity between generated and reference texts. This enables a more reliable evaluation of whether the meaning of a response is preserved, even when different phrasings are used. However, while BERTScore provides strong advantages in measuring semantic consistency, it is less suited for scenarios involving long-form, explanatory, or unanswerable responses. To address this gap, Generated Accuracy (G-Acc) (Ma et al., 2024c) has been proposed, which extends evaluation to free-form answers that emphasize reasoning, elaboration, and contextual completeness, thereby offering a more comprehensive assessment of generation quality.